OpenStudy (anonymous):

if 10^50-74 is written as an integer in base 10 notation wat is the sum of the digits in that integer

OpenStudy (anonymous):

ans is 440 but how do v solve tis??

OpenStudy (anonymous):

(10^4) - 74 = ? (10^4) - 10*7.4 = ? 10*[(10^3) - 7.4] = ? 10*[10^3 - 7.4] = 992.6*10 = 9926 = 9*2+2+6 = 26 Note that 10^4 will yield two nines a six and a two. so solving for 10^50 and 74 will give 48 nines a six and a two: 10*[10^49 - 7.4] = 9*48+2+6 = 440

OpenStudy (anonymous):

@Fall12-13 wer did u get tis frm i hav the exact soln wid me but didnt get it as in how 10^50 gives 48 nines

OpenStudy (anonymous):

Can you post your steps/how you worked the problem? We can see where you're going wrong then.

OpenStudy (anonymous):

solve tis 10(10^49-7.4)

OpenStudy (anonymous):

You get 48 nines,but the LAST numbers are 2 and 6, we take them in and add them. 10(10^49-7.4)=99999999999999999999999999999999999999999999999926

OpenStudy (anonymous):

got it... it was silly ques!! thanks:)

OpenStudy (anonymous):

No problem, nice to know.